# Quantum Molecular Charge-Transfer Model for Multi-step Auger-Meitner Decay Cascade Dynamics

### v1

@@ III.2 Decay Spawning Dynamics
Fig. 6 demonstrates dynamics of a sequence of 20 successive spawning events in both Channel 1 and 2, resulting in 20 trajectories in the dication states and 191 trajectories in the trication states in each channel. Fig. 6 a) shows the change in bond lengths as new trajectories are spawned. The solid black line is the initial IBr[]+1 state. At each spawning event a single new trajectory is spawned on the dication AIMD surface (dark green solid lines for Channel 1, dark purple solid lines for Channel 2) at the point of closest bond length to the IBr[]+1 state at . The spawning of trajectories determined by maximum structural similarity is then mapped to other molecular properties from the AIMD simulation. Fig. 6 b) shows the kinetic energies of the spawning decay trajectories. Unlike the bond lengths, there is now a discontinuous evolution in the kinetic energy between different surfaces as a result of the trajectories spawning on potential energy surfaces with steeper energy gradients. Fig. 6 c) and d) show the Br and I atomic charges respectively, they also have a discontinuous evolution as the total charge increases by 1 between the spawning events.
To give the distribution of trajectories shown in Fig. 6 physical meaning, we assign population weights from the expanding basis of quantum amplitudes trajectories from the spawning decay dynamics simulation. Fig. 7 a) and b) show the evolution of the state populations over time for the individual amplitude trajectories and for the sum of these trajectories respectively. Between and the first spawning decay event (), the population of the initial IBr[]+1 state is depleted by the Auger-Meitner decay process to the two dictation states. At , the population loss from the IBr[]+1 state is transferred to the dictation states, renormalizing the total population of the wavepacket to 1. Between and , the population of the dication states diminishes, and at , this population loss is transferred to the trication states. The population transfer from dication to trication states continues for the remainder of the simulation.
From the ratio of the population of IBr[4s1,V2]+3 to IBr[4s2,V1]+3 at 242 fs shown in Fig. 7, our model shows branching ratio of and fragmentation channels to be about 3. This calculated ratio is slightly higher than the experimental value. This overestimation can be due to our simplified model, which includes only 5 states.
@@
The evolution of the total population-weighted atomic charge and spin provides a quantum molecular charge transfer description of the multistep, multichannel Auger decay cascade. Although a more frequent spawning recurrence would create a smoother charge transfer curve, the recurrence interval used in this simulation effectively captures the charge flow between the two atoms during the coupled Auger-Meitner decay and fragmentation dynamics. In the first 50 fs, there is a flow of charge towards the Br atom, followed by a charge flow towards the iodine atom. This charging process converges around 75 fs, which can be approximated as the total charge transfer time for this decay cascade model.
@@ IV Conclusion
To demonstrate the construction of the decay spawning algorithm and dynamics simulation, this study uses a simplified decay cascade model for IBr with a core hole in the Br 3p orbital and focuses on the subsequent fragmentation dynamics that generate ion pairs of and . Our model predicts the ratios of these two ion pairs and their kinetic energy that agree with the experimental values. Our current implementation uses a single non-interacting potential energy surface for one possible electronic configuration of each state and restricts the configuration space in the electronic structure calculation to ensure the production of smooth potential energy surfaces. In future implementations, the decay spawning algorithm can be coupled with ”on-the-fly” trajectory surface hopping or AIMS simulations. This could enable modelling both the Auger-Meitner decay across multiple charge states and the non-adiabatic dynamics occurring between potential energy surfaces of the same charge state. This would provide a powerful exploratory description of the fragmentation dynamics following inner-shell decay processes in heavy element-containing molecules.

### v2

@@ III.2 Decay Spawning Dynamics
Fig. 6 demonstrates dynamics of a sequence of 20 successive spawning events in both Channel 1 and 2, resulting in 20 trajectories in the dication states and 191 trajectories in the trication states in each channel. Fig. 6 a) shows the change in bond lengths as new trajectories are spawned. The solid black line is the initial IBr[]+1 state. At each spawning event, a single new trajectory is spawned on the dication AIMD surface (dark green solid lines for Channel 1, dark purple solid lines for Channel 2) at the point of closest bond length to the IBr[]+1 state at . The spawning of trajectories determined by maximum structural similarity is then mapped to other molecular properties from the AIMD simulation. Fig. 6 b) shows the kinetic energies of the spawning decay trajectories. Unlike the bond lengths, there is now a discontinuous evolution in the kinetic energy between different surfaces as a result of the trajectories spawning on potential energy surfaces with steeper energy gradients. Fig. 6 c) and d) show the Br and I atomic charges respectively, they also have a discontinuous evolution as the total charge increases by 1 between the spawning events.
To give the distribution of trajectories shown in Fig. 6 physical meaning, we assign population weights from the expanding basis of quantum amplitudes trajectories from the spawning decay dynamics simulation. Fig. 7 a) and b) show the evolution of the state populations over time for the individual amplitude trajectories and for the sum of these trajectories respectively. Between and the first spawning decay event (), the population of the initial IBr[]+1 state is depleted by the Auger-Meitner decay process to the two dictation states. At , the population loss from the IBr[]+1 state is transferred to the dictation states, renormalizing the total population of the wavepacket to 1. Between and , the population of the dication states is reduced, and at , this population loss is transferred to the trication states. The population transfer from dication to trication states continues for the remainder of the simulation.
From the ratio of the population of IBr[,,]+3 to IBr[,]+3 at 242 fs shown in Fig. 7, our model shows branching ratio of and fragmentation channels to be about 3. This calculated ratio is slightly higher than the experimental value of 1.5. This overestimation can be due to our simplified model, which includes only 5 states.
@@
The evolution of the total population-weighted atomic charge in Fig. 8 c) provides a quantum molecular charge transfer description of the multistep, multichannel Auger decay cascade. Although a more frequent spawning recurrence would create a smoother charge transfer curve, the recurrence interval used in this simulation effectively captures the charge flow between the two atoms during the coupled Auger-Meitner decay and fragmentation dynamics. In the first 50 fs, the charge of IBr increases due to Auger-Meitner decays, followed by a charge flow towards the iodine atom. This process converges around 75 fs, which can be approximated as the total charge transfer time for this decay cascade model.
@@ IV Conclusion
To demonstrate our the decay spawning algorithm and dynamics simulation, this study uses a simplified decay cascade model for IBr with a core hole in the Br 3p orbital and focuses on the subsequent fragmentation dynamics that generate ion pairs of and . Our model predicts the ratios of these two ion pairs and their kinetic energy that agree with the experimental values. Our current implementation uses a single non-interacting potential energy surface for one possible electronic configuration of each state and restricts the configuration space in the electronic structure calculation to ensure the production of smooth potential energy surfaces. In future implementations, the decay spawning algorithm can be coupled with “on-the-fly” trajectory surface hopping[39] or multiple cloning (MC) methods[40]. Multiple cloning dynamics methods combine the time expanding basis in AIMS with the propagation of an average trajectory used in multiconfigurational Ehrenfest dynamics[41] and have been recently developed to treat a dense manifold of electronic states[42]. This approach could potentially enable modeling both the Auger-Meitner decay across multiple charge states and the non-adiabatic dynamics occurring between potential energy surfaces of the same charge state. Furthermore, by including more orbitals and decay steps, our approach would provide a powerful exploratory description of the fragmentation dynamics following inner-shell decay processes in heavy element-containing molecules.